# refrences.bib

### v1

@@ Conclusion
    From Theorem 3, $S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)=\bigcup_{e^{\prime}\in E^{\prime}}S\left(e^{\prime}\right)$, where $E^{\prime}$ is the set of all the restrictions of admissible functions $e\in E$ to $\mathscr{I}-\left\{i_{0}\right\}$. So, for each $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, there exists at least one $e_{0}^{\prime}\in E^{\prime}$ such that $x\in S\left(e_{0}^{\prime}\right)$. Now, from (11), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}$ or $x_{j}\in I_{j}$. But, in the former case, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 4), we have again $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}\subseteq I_{j}$. Consequently, $x_{j}\in I_{j}$ (and therefore $x_{j}\in I_{i_{0}j}$ from Definition 4), $\forall j\in\mathscr{J}$. ∎
  • The proof of Lemma 7.
[…]
  • The proof of Lemma 8.
[…]
  • The proof of Lemma 9.
@@
    From $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$ and Lemma 6, we conclude that $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}\left(*1\right)$. Also, from Lemma 5, there exists at least one $j_{i}\in\mathscr{J}$ such that $x_{j_{i}}\in S_{ij_{i}}^{\prime}$. But, since $S_{ij}^{\prime}\subseteq S_{i_{0}j}^{\prime}$, it follows that $x_{j_{i}}\in S_{i_{0}j_{i}}^{\prime}(*2)$. So, (1),(2)(*1),(*2)( ∗ 1 ) , ( ∗ 2 ) and Corollary 5 imply that $x$ satisfies the $i_{0}$’th equation. ∎
  • The proof of Lemma 10.
@@
    If $x\in S\left(A^{+},A^{-},b\right)$, then from Lemma 5, there exists at least one $j\in\mathscr{J}$ such that $x_{j}\in S_{i_{0}j}^{\prime}$. But, since $\mathscr{J}_{i_{0}}=\left\{j_{0}\right\}$, we have necessarily $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}=\{k\}$, that means $x_{j_{0}}=k$. Moreover, if $x\in S_{\{i\}}\left(A^{+},A^{-},b\right)\cap\left\{x\in[0,1]^{n}:x_{j_{0}}=k\right\}$, then $x_{j_{0}}=k\in S_{ij_{0}}^{\prime}$ and also from Lemma 6, $x_{j}\in I_{ij}$, $\forall j\in\mathscr{J}$. So, Corollary 5 implies that $x$ satisfies the $i$’th equation. ∎
  • The proof of Lemma 11.
[…]
  • The proof of Lemma 12.
@@
    Let $e\in E$. So, from (15), we have either $x(e)_{j_{0}}=L_{j_{0}}$ or $x(e)_{j_{0}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}$. In the latter case, it is concluded that $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{\mathrm{ij}_{0}}^{\prime}\neq\varnothing$ (Corollary 6), which requires $\mathscr{I}_{j_{0}}(e)\subseteq\mathscr{I}_{j_{0}}$. Hence, by our assumption, we have $L_{j_{0}}\in S_{ij_{0}}^{\prime}$, $\forall i\in\mathscr{I}_{j_{0}}(e)$, that implies $\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}=L_{j_{0}}$. So, in any case, $x(e)_{j_{0}}=L_{j_{0}}$. Now, let $L_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$ and $x(e)\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. So, from Lemma 6, it follows that $x(e)_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}~{}(*)$. Moreover, we have either $x(e)_{j_{0}}=L_{j_{0}}$ or $x(e)_{j_{0}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}=L_{j_{0}}$, where the latter equality results from the assumption that $L_{j_{0}}\in S_{ij_{0}}^{\prime}$, $\forall i\in\mathscr{I}_{j_{0}}$. Hence, $x(e)_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$, which together with ()(*)( ∗ ) and Corollary 5 imply that $x(e)$ satisfies the $i_{0}$’th equation. ∎
  • The proof of Lemma 13.
@@
[…]

### v2

@@ Conclusion
From Theorem 3.10, $S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)=\bigcup_{e^{\prime}\in E^{\prime}}S\left(e^{\prime}\right)$, where $E^{\prime}$ is the set of all the restrictions of admissible functions $e\in E$ to $\mathscr{I}-\left\{i_{0}\right\}$. So, for each $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, there exists at least one $e_{0}^{\prime}\in E^{\prime}$ such that $x\in S\left(e_{0}^{\prime}\right)$. Now, from (11), for each $j\in\mathscr{J}$ we have either $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}$ or $x_{j}\in I_{j}$. But, in the former case, since $S_{ij}^{\prime}=S_{ij}\cap I_{j}$ (Definition 2.15), we have again $x_{j}\in\bigcap_{i\in\mathscr{I}_{j}\left(e_{0}^{\prime}\right)}S_{ij}^{\prime}\subseteq I_{j}$. Consequently, $x_{j}\in I_{j}$ (and therefore $x_{j}\in I_{i_{0}j}$ from Definition 2.15), $\forall j\in\mathscr{J}$. ∎
The proof of Lemma 5.2.
@@
To prove the lemma, it is sufficient to show that $S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)\subseteq S\left(A^{+},A^{-},b\right)$. In other words, we shall show that if $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, then $x$ satisfies the $i_{0}$’th equation. For this purpose, we first note from Lemma 5.1 that $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$ $\left(*1\right)$ and particularly $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}$ $\left(*2\right)$. Also, since $S\left(A^{+},A^{-},b\right)\neq\varnothing$, there exists at least one $j_{0}\in\mathscr{J}$ such that $S_{i_{0}j_{0}}^{\prime}\neq\varnothing$ (Lemma 3.1 (b)). So, from $b_{i_{0}}=0$ and Remark 2.13, we have $S_{i_{0}j_{0}}^{\prime}=\left[1-u_{i_{0}j_{0}}^{-},u_{i_{0}j_{0}}^{+}\right]$ which together with Corollary 2.16 imply $S_{i_{0}j_{0}}^{\prime}=\left[L_{j_{0}},U_{j_{0}}\right]=I_{j_{0}}(*3)$. Hence, from $\left(*1\right)$ and $\left(*3\right)$, it follows that $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}\left(*4\right)$. Now, the result follows from (2),(4)(*2),(*4) and Corollary 3.4. ∎
The proof of Lemma 5.4.
@@
For each $x\in S\left(A^{+},A^{-},b\right)$, Lemma 3.2 implies that $x_{j_{0}}\in I_{j_{0}}=\{k\}$, i.e., $x_{j_{0}}=k$. Now, from Lemma 5.1, if $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$, then we have $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}\left(*1\right)$. Particularly, $x_{j_{0}}\in I_{i_{0}j_{0}}\subseteq I_{j_{0}}=\{k\}$. On the other hand, since $k\in S_{i_{0}{j_{0}}}^{\prime}$, then we have $S_{i_{0}j_{0}}^{\prime}\neq\varnothing$, that together with $S_{i_{0}j_{0}}^{\prime}=S_{i_{0}j_{0}}\cap I_{j_{0}}$ and $I_{j_{0}}=\{k\}$ imply $S_{i_{0}j_{0}}^{\prime}=\{k\}$. Consequently, $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}\left(*2\right)$. Now, $\left(*1\right)$$\left(*2\right)$ and Corollary 3.4 imply that $x$ satisfies the $i_{0}$’th equation. ∎
The proof of Lemma 5.6.
From $x\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$ and Lemma 5.1, we conclude that $x_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}\left(*1\right)$. Also, from Lemma 3.2, there exists at least one $j_{i}\in\mathscr{J}$ such that $x_{j_{i}}\in S_{ij_{i}}^{\prime}$. But, since $S_{ij}^{\prime}\subseteq S_{i_{0}j}^{\prime}$, it follows that $x_{j_{i}}\in S_{i_{0}j_{i}}^{\prime}(*2)$. So, (1),(2)(*1),(*2) and Corollary 3.4 imply that $x$ satisfies the $i_{0}$’th equation. ∎
The proof of Lemma 5.8.
If $x\in S\left(A^{+},A^{-},b\right)$, then from Lemma 3.2, there exists at least one $j\in\mathscr{J}$ such that $x_{j}\in S_{i_{0}j}^{\prime}$. But, since $\mathscr{J}_{i_{0}}=\left\{j_{0}\right\}$, we have necessarily $x_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}=\{k\}$, that means $x_{j_{0}}=k$. Moreover, if $x\in S_{\{i\}}\left(A^{+},A^{-},b\right)\cap\left\{x\in[0,1]^{n}:x_{j_{0}}=k\right\}$, then $x_{j_{0}}=k\in S_{ij_{0}}^{\prime}$ and also from Lemma 5.1, $x_{j}\in I_{ij}$, $\forall j\in\mathscr{J}$. So, Corollary 3.4 implies that $x$ satisfies the $i$’th equation. ∎
The proof of Lemma 5.10.
@@
Assume that $x(e)\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. To prove the lemma, it is sufficient to show that $x(e)$ also satisfies the $i_{0}$’th equation. First, from Lemma 5.1, it follows that $x(e)_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}\left(*1\right)$. On the other hand, from $\left|S_{i_{0}j_{0}}^{\prime}\right|=2$ and Corollary 2.16, we have $S_{i_{0}j_{0}}^{\prime}=\left\{L_{j_{0}},U_{j_{0}}\right\}$. Moreover, based on (12), it follows that either $x(e)_{j_{0}}=L_{j_{0}}$ (if $\mathscr{I}_{j_{0}}(e)=\varnothing$) or $x(e)_{j_{0}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}$ (if $\mathscr{I}_{j_{0}}(e)\neq\varnothing$ ). In the former case, it is clear that $x(e)_{j_{0}}=L_{j_{0}}\in S_{i_{0}j_{0}}$. Otherwise, in the latter case, it follows that $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\neq\varnothing$ (Corollary 3.8) and $\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}\in\left\{L_{j_{0}},U_{j_{0}}\right\}$ (Corollary 2.16), and therefore in any case we have $x(e)_{j_{0}}\in S_{i_{0}j_{0}}\left(*2\right)$. As a result, $\left(*1\right)$, $\left(*2\right)$ and Corollary 3.4 imply that $x(e)$ also satisfies the $i_{0}$’th equation. ∎
The proof of Lemma 5.13.
Let $e\in E$. So, from (12), we have either $x(e)_{j_{0}}=L_{j_{0}}$ or $x(e)_{j_{0}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}$. In the latter case, it is concluded that $\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{\mathrm{ij}_{0}}^{\prime}\neq\varnothing$ (Corollary 3.8), which requires $\mathscr{I}_{j_{0}}(e)\subseteq\mathscr{I}_{j_{0}}$. Hence, by our assumption, we have $L_{j_{0}}\in S_{ij_{0}}^{\prime}$, $\forall i\in\mathscr{I}_{j_{0}}(e)$, that implies $\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}=L_{j_{0}}$. So, in any case, $x(e)_{j_{0}}=L_{j_{0}}$. Now, let $L_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$ and $x(e)\in S_{\left\{i_{0}\right\}}\left(A^{+},A^{-},b\right)$. So, from Lemma 5.1, it follows that $x(e)_{j}\in I_{i_{0}j}$, $\forall j\in\mathscr{J}~(*)$. Moreover, we have either $x(e)_{j_{0}}=L_{j_{0}}$ or $x(e)_{j_{0}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{0}}(e)}S_{ij_{0}}^{\prime}\right\}=L_{j_{0}}$, where the latter equality results from the assumption that $L_{j_{0}}\in S_{ij_{0}}^{\prime}$, $\forall i\in\mathscr{I}_{j_{0}}$. Hence, $x(e)_{j_{0}}\in S_{i_{0}j_{0}}^{\prime}$, which together with ()(*) and Corollary 3.4 imply that $x(e)$ satisfies the $i_{0}$’th equation. ∎
The proof of Lemma 5.15.
@@
Hence, $\mathscr{I}_{j_{1}}\left(e_{0}\right)=\varnothing$ that implies $e\in E_{0}$, $\mathscr{I}_{j_{1}}\left(e_{0}\right)\subseteq\mathscr{I}_{j_{1}}(e)$ and $x\left(e_{0}\right)_{j_{1}}=L_{j_{1}}$ (from (12)). However, $\bigcap_{i\in\mathscr{I}_{j_{1}}(e)}S_{ij_{1}}^{\prime}\neq\varnothing$ (Corollary 3.8) and $\min\left\{\bigcap_{i\in\mathscr{I}_{j_{1}}(e)}S_{ij_{1}}^{\prime}\right\}\in\left\{L_{j_{1}},U_{j_{1}}\right\}$ (Corollary 2.16), which requires $x(e)_{j_{1}}=v\in\left\{L_{j_{1}},U_{j_{1}}\right\}$. So, $c_{j_{1}}x\left(e_{0}\right)_{j_{1}}=c_{j_{1}}L_{j_{1}}\leq c_{j_{1}}v=c_{j_{1}}x(e)_{j_{1}}\left(*1\right)$. Moreover, based on (16), we have $\mathscr{I}_{j}\left(e_{0}\right)=\mathscr{I}_{j}(e)$, $\forall j\in\mathscr{J}-\left\{j_{1},j_{2}\right\}$, which means $x\left(e_{0}\right)_{j}=x(e)_{j}$, and therefore $c_{j}x\left(e_{0}\right)_{j}=c_{j}x(e)_{j}$, $\forall j\in\mathscr{J}-\left\{j_{1},j_{2}\right\}(*2)$. Subsequently, the assumption $\bigcap_{i\in\mathscr{I}_{j_{2}}}S_{ij_{2}}^{\prime}=\left\{L_{j_{2}}\right\}$ together with Remark 5.12 imply that $x\left(e_{0}\right)_{j_{2}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{2}}\left(e_{0}\right)}S_{ij_{2}}^{\prime}\right\}=L_{j_{2}}$. Also, if $\mathscr{I}_{j_{2}}(e)=\varnothing$ we have $x(e)_{j_{2}}=L_{j_{2}}$ (from (12)) and if $\mathscr{I}_{j_{2}}(e)\neq\varnothing$ we have $x(e)_{j_{2}}=\min\left\{\bigcap_{i\in\mathscr{I}_{j_{2}}(e)}S_{ij_{2}}^{\prime}\right\}=L_{j_{2}}$ (Remark 5.12 and (12)). Hence, $x\left(e_{0}\right)_{j_{2}}=x(e)_{j_{2}}$ that requires $c_{j_{2}}x\left(e_{0}\right)_{j_{2}}=c_{j_{2}}x(e)_{j_{2}}\left(*3\right)$. Now, from $\left(*1\right)$, $\left(*2\right)$ and (3)(*3), $c_{j_{1}}x\left(e_{0}\right)_{j_{1}}+c_{j_{2}}x\left(e_{0}\right)_{j_{2}}+\sum_{j\in\mathscr{J}-\left\{j_{1},j_{2}\right\}}c_{j}x\left(e_{0}\right)_{j}\leq c_{j_{1}}x(e)_{j_{1}}+c_{j_{2}}x(e)_{j_{2}}+\sum_{j\in\mathscr{J}-\left\{j_{1},j_{2}\right\}}c_{j}x(e)_{j}$. (b) It is quite similar to the proof of Part (a). ∎
Acknowledgement
The authors wish to express their appreciation for several excellent suggestions for improvements in this paper made by the referees.